A concave cosmetic mirror has a focal length of 50
Answer #1

ho = height of object (brush)
f = focal length
do = dist of object
di = dist of image
hi = height of image

ho = 4 cm do = 25 cm f = 50 cm

di = ? hi = ?

To determine the image distance, use the mirror equation

1/f = 1/do + 1/di
1/(50 cm) = 1/(25 cm) + 1/di

di = -50 cm

Height of image:

hi/ho = - di/do
hi /(4.0 cm) = - (-50 cm)/(25 cm)

hi = 8 cm